# X-Dreamer: Creating High-quality 3D Content by Bridging the Domain Gap Between Text-to-2D and Text-to-3D Generation

### v2

@@
Abstract
In recent times, automatic text-to-3D content creation has made significant progress, driven by the development of pretrained 2D diffusion models. Existing text-to-3D methods typically optimize the 3D representation to ensure that the rendered image aligns well with the given text, as evaluated by the pretrained 2D diffusion model. Nevertheless, a substantial domain gap exists between 2D images and 3D assets, primarily attributed to variations in camera-related attributes and the exclusive presence of foreground objects. Consequently, employing 2D diffusion models directly for optimizing 3D representations may lead to suboptimal outcomes. To address this issue, we present X-Dreamer, a novel approach for high-quality text-to-3D content creation that effectively bridges the gap between text-to-2D and text-to-3D synthesis. The key components of X-Dreamer are two innovative designs: Camera-Guided Low-Rank Adaptation (CG-LoRA) and Attention-Mask Alignment (AMA) Loss. CG-LoRA dynamically incorporates camera information into the pretrained diffusion models by employing camera-dependent generation for trainable parameters. This integration enhances the alignment between the generated 3D assets and the camera’s perspective. AMA loss guides the attention map of the pretrained diffusion model using the binary mask of the 3D object, prioritizing the creation of the foreground object. This module ensures that the model focuses on generating accurate and detailed foreground objects. Extensive evaluations demonstrate the effectiveness of our proposed method compared to existing text-to-3D approaches. Our project webpage: https://xmu-xiaoma666.github.io/Projects/X-Dreamer.
1 Introduction
@@
Figure 1: The outputs of the text-to-2D generation model (left) and the text-to-3D generation model (right) under the same text prompt, i.e., “A statue of Leonardo DiCaprio’s head”.
@@
In recent years, extensive research has demonstrated significant performance improvement in the text-to-2D generation task [24, 26, 25, 2] by leveraging pretrained diffusion models [30, 6, 31] on a large-scale text-image dataset [27]. Building on these advancements, DreamFusion [22] introduces an effective approach that utilizes a pretrained 2D diffusion model [26] to autonomously generate 3D assets from text, eliminating the need for a dedicated 3D asset dataset. A key innovation introduced by DreamFusion is the Score Distillation Sampling (SDS) algorithm. This algorithm aims to optimize a single 3D representation, such as NeRF [18], to ensure that rendered images from any camera perspective maintain a high likelihood with the given text, as evaluated by the pretrained 2D diffusion model. Inspired by the groundbreaking SDS algorithm, several recent works [13, 4, 16, 35, 36] have emerged, envisioning the text-to-3D generation task through the application of pretrained 2D diffusion models.
@@
While text-to-3D generation has made significant strides through the utilization of pretrained text-to-2D diffusion models, it is crucial to recognize and address the persistent and substantial domain gap that remains between text-to-2D and text-to-3D generation. This distinction is clearly illustrated in Fig. 1. To begin with, the text-to-2D model produces camera-independent generation results, focusing on generating high-quality images from specific angles while disregarding other angles. In contrast, 3D content creation is intricately tied to camera parameters such as position, shooting angle, and field of view. As a result, a text-to-3D model must generate high-quality results across all possible camera parameters. This fundamental difference emphasizes the necessity for innovative approaches that enable the pretrained diffusion model to consider camera parameters. Furthermore, a text-to-2D generation model must simultaneously generate both foreground and background elements while maintaining the overall coherence of the image. Conversely, a text-to-3D generation model only needs to concentrate on creating the foreground object. This distinction allows text-to-3D models to allocate more resources and attention to precisely represent and generate the foreground object. Consequently, the domain gap between text-to-2D and text-to-3D generation poses a significant performance obstacle when directly employing pretrained 2D diffusion models for 3D asset creation.
@@
In this study, we present a pioneering framework, X-Dreamer, designed to address the domain gap between text-to-2D and text-to-3D generation, thereby facilitating the creation of high-quality text-to-3D content. Our framework incorporates two innovative designs that are specifically tailored to address the aforementioned challenges. Firstly, existing approaches [13, 4, 16, 35] commonly employ 2D pretrained diffusion models [26, 25] for text-to-3D generation, which lack inherent linkage to camera parameters. To address this limitation and ensure that our text-to-3D model produces results that are directly influenced by camera parameters, we introduce Camera-Guided Low-Rank Adaptation (CG-LoRA) to fine-tune the pretrained 2D diffusion model. Notably, the parameters of CG-LoRA are dynamically generated based on the camera information during each iteration, establishing a robust relationship between the text-to-3D model and camera parameters. Furthermore, pretrained text-to-2D diffusion models allocate attention to both foreground and background generation, whereas the creation of 3D assets necessitates a stronger focus on accurately generating foreground objects. To address this requirement, we introduce Attention-Mask Alignment (AMA) Loss, which leverages the rendered binary mask of the 3D object to guide the attention map of the pretrained 2D stable diffusion model [25]. By incorporating this module, X-Dreamer prioritizes the generation of foreground objects, resulting in a significant enhancement of the overall quality of the generated 3D content.
@@
We present a compelling demonstration of the effectiveness of X-Dreamer in synthesizing high-quality 3D assets based on textual cues. By incorporating CG-LoRA and AMA loss to address the domain gap between text-to-2D and text-to-3D generation, our proposed framework exhibits substantial advancements over prior methods in text-to-3D generation. In summary, our study contributes to the field in three key aspects:
@@
  • To enhance the alignment between the generated results and the camera perspective, we propose CG-LoRA, which leverages camera information to dynamically generate specific parameters for 2D diffusion models.
@@
2 Related Work
Figure 2: Overview of the proposed X-Dreamer, which consists of geometry learning and appearance learning.
2.1 Text-to-3D Content Creation
@@
In recent years, there has been a significant surge in interest surrounding the evolution of text-to-3D generation [17, 22, 12]. This growing field has been propelled, in part, by advancements in pretrained vision-and-language models, such as CLIP [23], as well as diffusion models like Stable Diffusion [25] and Imagen [26]. Contemporary text-to-3D models can generally be classified into two distinct categories: the CLIP-based text-to-3D approach and the diffusion-based text-to-3D approach. The CLIP-based text-to-3D approach [17, 14, 11, 9, 19, 37] employs CLIP encoders [23] to project textual descriptions and rendered images derived from the 3D object into a modal-shared feature space. Subsequently, CLIP loss is harnessed to align features from both modalities, optimizing the 3D representation to conform to the textual description. Various scholars have made significant contributions to this field. For instance, Michel et al. [17] are pioneers in proposing the use of CLIP loss to harmonize the text prompt with the rendered images of the 3D object, thereby enhancing text-to-3D generation. Ma et al. [14] introduce dynamic textual guidance during 3D object synthesis to improve convergence speed and generation performance. However, these approaches have inherent limitations, as they tend to generate 3D representations with a scarcity of geometry and appearance detail. To overcome this shortcoming, the diffusion-based text-to-3D approach [22, 13, 4, 32, 10, 8] leverages pretrained text-to-2D diffusion models [25, 26] to guide the optimization of 3D representations. Central to these models is the application of SDS loss [22] to align the rendered images stemming from a variety of camera perspectives with the textual description. Specifically, given the target text prompt, Lin et al. [13] leverage a coarse-to-fine pipeline to generate high-resolution 3D content. Chen et al. [4] decouple geometry modeling and appearance modeling to generate realistic 3D assets. For specific purposes, some researchers [28, 36] integrate trainable LoRA [7] branches into pretrained diffusion models. For instance, Seo et al. [28] put forth 3DFuse, a model that harnesses the power of LoRA to comprehend object semantics. Wang et al. [36] introduce ProlificDreamer, where the role of LoRA is to evaluate the score of the variational distribution for 3D parameters. However, the LoRA parameter begins its journey from random initialization and maintains its independence from the camera and text. To address these limitations, we present two innovative modules: CG-LoRA and AMA loss. These modules are designed to enhance the model’s ability to consider important camera parameters and prioritize the generation of foreground objects throughout the text-to-3D creation process.
2.2 Low-Rank Adaptation (LoRA)
@@
Low-Rank Adaptation (LoRA) is a technique used to reduce memory requirements when fine-tuning a large model. It involves injecting only a small set of trainable parameters into the pretrained model, while keeping the original parameters fixed. During the optimization process, gradients are passed through the fixed pretrained model weights to the LoRA adapter, which is then updated to optimize the loss function. LoRA has been applied in various fields, including natural language processing [7, 3], image synthesis [38] and 3D generation [28, 36]. To achieve low-rank adaptation, a linear projection with a pretrained weight matrix $\mathbf{W}_{0}\in\mathbb{R}^{d_{in}\times d_{out}}$ is augmented with an additional low-rank factorized projection. This augmentation is represented as $\mathbf{W}_{0}+\Delta\mathbf{W}=\mathbf{W}_{0}+\mathbf{A}\mathbf{B}$, where $\mathbf{A}\in\mathbb{R}^{d_{in}\times r}$, $\mathbf{B}\in\mathbb{R}^{r\times d_{out}}$, and $r\ll\min(d_{in},d_{out})$. During training, $\mathbf{W}_{0}$ remains fixed, while $\mathbf{A}$ and $\mathbf{B}$ are trainable. The modified forward pass, given the original forward pass $\mathbf{Y}=\mathbf{X}\mathbf{W}_{0}$, can be formulated as follows:
$\mathbf{Y}=\mathbf{X}\mathbf{W}_{0}+\mathbf{X}\mathbf{A}\mathbf{B}.$ (1)
@@
3 Approach
Figure 3: Illustration of Camera-Guided Low-Rank Adaptation.
3.1 Architecture
@@
In this section, we present a comprehensive introduction to the proposed X-Dreamer, which consists of two main stages: geometry learning and appearance learning. For geometry learning, we employ DMTet [29] as the 3D representation. DMTet is an MLP parameterized with $\Phi_{dmt}$ and is initialized with a 3D ellipsoid using the mean squared error (MSE) loss $\mathcal{L}_{MSE}$. Subsequently, we optimize DMTet and CG-LoRA using the SDS loss [22] $\mathcal{L}_{SDS}$ and the proposed AMA loss $\mathcal{L}_{AMA}$ to ensure the alignment between the 3D representation and the input text prompt. For appearance learning, we leverage bidirectional reflectance distribution function (BRDF) modeling [33] following the previous approach [4]. Specifically, we utilize an MLP with trainable parameters $\Phi_{mat}$ to predict surface materials. Similar to the geometry learning stage, we optimize $\Phi_{mat}$ and CG-LoRA using the SDS loss $\mathcal{L}_{SDS}$ and the AMA loss $\mathcal{L}_{AMA}$ to achieve alignment between the 3D representation and the text prompt. Fig. 2 provides a detailed depiction of our proposed X-Dreamer.
@@
Geometry Learning. For geometry learning, an MLP network $\Phi_{dmt}$ is utilized to parameterize DMTet as a 3D representation. To enhance the stability of geometry modeling, we employ a 3D ellipsoid as the initial configuration for DMTet $\Phi_{dmt}$. For each vertex $v_{i}\in V_{T}$ belonging to the tetrahedral grid $T$, we train $\Phi_{dmt}$ to predict two important values: the SDF value $s(v_{i})$ and the deformation offset $\delta(v_{i})$. To initialize $\Phi_{dmt}$ with the 3D ellipsoid, we sample a set of $N$ points $\{p_{i}\in\mathbb{R}^{3}\}|_{i=1}^{N}$ approximately distributed on the surface of an ellipsoid and compute the corresponding SDF values $\{{SDF}(p_{i})\}|_{i=1}^{N}$. Subsequently, we optimize $\Phi_{dmt}$ using MSE loss. This optimization process ensures that $\Phi_{dmt}$ effectively initializes DMTet to resemble the 3D ellipsoid. The formulation of the MSE loss is given by:
$\mathcal{L}_{MSE}=\frac{1}{N}\sum_{i=1}^{N}\left(s(p_{i};\Phi_{dmt})-SDF(p_{i})\right)^{2}.$ (2)
@@
After initializing the geometry, our objective is to align the geometry of DMTet with the input text prompt. Specifically, we generate the normal map $\boldsymbol{n}$ and the object mask $\boldsymbol{m}$ from the initialized DMTet $\Phi_{dmt}$ by employing a differentiable rendering technique [33], given a randomly sampled camera pose $\boldsymbol{c}$. Subsequently, we input the normal map $\boldsymbol{n}$ into the frozen stable diffusion (SD) with a trainable CG-LoRA and update $\Phi_{dmt}$ using the SDS loss, which is defined as follows:
@@
$\small\nabla_{\Phi_{dmt}}\mathcal{L}_{\mathrm{SDS}}=\mathbb{E}_{t,\boldsymbol{\epsilon}}\left[w(t)\left(\hat{\boldsymbol{\epsilon}}_{\Theta}(\boldsymbol{n}_{t};\boldsymbol{y},t)-\boldsymbol{\epsilon}\right)\frac{\partial\boldsymbol{n}}{\partial{\Phi_{dmt}}}\right],$ (3)
@@
where $\Theta$ represents the parameter of SD, $\hat{\boldsymbol{\epsilon}}_{\Theta}(\boldsymbol{n}_{t};\boldsymbol{y},t)$ denotes the predicted noise of SD given the noise level $t$ and text embedding $\boldsymbol{y}$. Additionally, $\boldsymbol{n}_{t}=\alpha_{t}\boldsymbol{n}+\sigma_{t}\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\boldsymbol{I})$ represents noise sampled from a normal distribution. The implementation of $w(t)$, $\alpha_{t}$, and $\sigma_{t}$ is based on the DreamFusion [22].
@@
$\mathcal{L}_{AMA}=\frac{1}{L}\sum_{i=1}^{L}|\boldsymbol{a}_{i}-\eta(\boldsymbol{m})|,$ (4)
@@
Appearance Learning. After obtaining the geometry of the 3D object, our objective is to compute its appearance using the Physically-Based Rendering (PBR) material model [15]. The material model comprises the diffuse term $\boldsymbol{k}_{d}\in\mathbb{R}^{3}$, the roughness and metallic term $\boldsymbol{k}_{rm}\in\mathbb{R}^{2}$, and the normal variation term $\boldsymbol{k}_{n}\in\mathbb{R}^{3}$. For any point $p\in\mathbb{R}^{3}$ on the surface of the geometry, we utilize an MLP parameterized by $\Phi_{mat}$ to obtain the three material terms, which can be expressed as follows:
$(\boldsymbol{k}_{d},\boldsymbol{k}_{n},\boldsymbol{k}_{rm})=\text{MLP}\left(\mathcal{P}(p);\Phi_{mat}\right),$ (5)

### v3

# Creating High-quality 3D Content by Bridging the Gap Between Text-to-2D and Text-to-3D Generation

@@
Abstract.
In recent times, automatic text-to-3D content creation has made significant progress, driven by the development of pretrained 2D diffusion models. Existing text-to-3D methods typically optimize the 3D representation to ensure that the rendered image aligns well with the given text, as evaluated by the pretrained 2D diffusion model. Nevertheless, a substantial domain gap exists between 2D images and 3D assets, primarily attributed to variations in camera-related attributes and the exclusive presence of foreground objects. Consequently, employing 2D diffusion models directly for optimizing 3D representations may lead to suboptimal outcomes. To address this issue, we present X-Dreamer, a novel approach for high-quality text-to-3D content creation that effectively bridges the gap between text-to-2D and text-to-3D synthesis. The key components of X-Dreamer are two innovative designs: Camera-Guided Low-Rank Adaptation (CG-LoRA) and Attention-Mask Alignment (AMA) Loss. CG-LoRA dynamically incorporates camera information into the pretrained diffusion models by employing camera-dependent generation for trainable parameters. This integration makes the 2D diffusion model camera-sensitive. AMA loss guides the attention map of the pretrained diffusion model using the binary mask of the 3D object, prioritizing the creation of the foreground object. This module ensures that the model focuses on generating accurate and detailed foreground objects. Extensive evaluations demonstrate the effectiveness of our proposed method compared to existing text-to-3D approaches. Our project webpage: https://anonymous-11111.github.io/. Our code is available at https://github.com/xmu-xiaoma666/X-Dreamer.
Text-to-3D Generation, Vision and Language, Domain Gap
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Applied computing Media artsccs: Applied computing Computer-aided designccs: Applied computing Computer-Fine arts
1. Introduction
Figure 1. The outputs of the text-to-2D generation model (left) and the text-to-3D generation model (right) under the same text prompt, i.e., “A statue of Leonardo DiCaprio’s head”.
Figure 2. Text-to-3D generation results of the proposed X-Dreamer.
@@ 1. Introduction
In recent years, extensive research (Chen et al., 2023b; Yu et al., 2023; Huang et al., 2023b) has demonstrated significant performance improvement in the text-to-2D generation task (Ramesh et al., 2022; Saharia et al., 2022; Rombach et al., 2022; Balaji et al., 2022; Wu et al., 2024) by leveraging pretrained diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020) on a large-scale text-image dataset (Schuhmann et al., 2022). Building on these advancements, DreamFusion (Poole et al., 2022) introduces an effective approach that utilizes a pretrained 2D diffusion model (Saharia et al., 2022) to autonomously generate 3D assets from text, eliminating the need for a dedicated 3D asset dataset. A key innovation introduced by DreamFusion is the Score Distillation Sampling (SDS) algorithm. This algorithm aims to optimize a single 3D representation, such as NeRF (Mildenhall et al., 2021), to ensure that rendered images from any camera perspective maintain a high likelihood with the given text, as evaluated by the pretrained 2D diffusion model. Inspired by the groundbreaking SDS algorithm, several recent works (Lin et al., 2023; Chen et al., 2023a; Metzer et al., 2023; Wang et al., 2023a, c) have emerged, envisioning the text-to-3D generation task through the application of pretrained 2D diffusion models.
@@
While text-to-3D generation has made significant strides through the utilization of pretrained text-to-2D diffusion models (Saharia et al., 2022; Zhang et al., 2023; Kumari et al., 2023), it is crucial to recognize and address the persistent and substantial domain gap that remains between text-to-2D and text-to-3D generation. This distinction is clearly illustrated in Fig. 1. To begin with, the text-to-2D model produces camera-independent generation results, focusing on generating high-quality images from specific angles while disregarding other angles. In contrast, 3D content creation is intricately tied to camera parameters such as position, shooting angle, and field of view. As a result, a text-to-3D model must generate high-quality results across all possible camera parameters. This fundamental difference emphasizes the necessity for innovative approaches that enable the pretrained diffusion model to consider camera parameters. Furthermore, a text-to-2D generation model (Sun et al., 2024; Chen et al., 2023c; Ruan et al., 2023; Qu et al., 2023; Lu et al., 2023; Fei et al., 2024a, c) must simultaneously generate both foreground and background elements while maintaining the overall coherence of the image. Conversely, a text-to-3D generation model (Raj et al., 2023; Xu et al., 2023; Lorraine et al., 2023) only needs to concentrate on creating the foreground object. This distinction allows text-to-3D models to allocate more resources and attention to precisely represent and generate the foreground object. Consequently, the domain gap between text-to-2D and text-to-3D generation poses a significant performance obstacle when directly employing pretrained 2D diffusion models for 3D asset creation.
@@
In this study, we present a pioneering framework, X-Dreamer, designed to address the domain gap between text-to-2D and text-to-3D generation, thereby facilitating the creation of high-quality text-to-3D content. Our framework incorporates two innovative designs that are specifically tailored to address the aforementioned challenges. Firstly, existing approaches (Lin et al., 2023; Chen et al., 2023a; Metzer et al., 2023; Wang et al., 2023a; Ma et al., 2022, 2023a) commonly employ 2D pretrained diffusion models (Saharia et al., 2022; Rombach et al., 2022) for text-to-3D generation, which lack inherent linkage to camera parameters. To address this limitation and ensure that our text-to-3D model produces results that are directly influenced by camera parameters, we introduce Camera-Guided Low-Rank Adaptation (CG-LoRA) to fine-tune the pretrained 2D diffusion model. Notably, the parameters of CG-LoRA are dynamically generated based on the camera information during each iteration, establishing a robust relationship between the text-to-3D model and camera parameters. Furthermore, pretrained text-to-2D diffusion models allocate attention to both foreground and background generation, whereas the creation of 3D assets necessitates a stronger focus on accurately generating foreground objects. To address this requirement, we introduce Attention-Mask Alignment (AMA) Loss, which leverages the rendered binary mask of the 3D object to guide the attention map of the pretrained 2D stable diffusion model (Rombach et al., 2022). By incorporating this module, X-Dreamer prioritizes the generation of foreground objects, resulting in a significant enhancement of the overall quality of the generated 3D content.
@@
We present a compelling demonstration of the effectiveness of X-Dreamer in synthesizing high-quality 3D assets based on textual cues, as shown in Fig. 2. By incorporating CG-LoRA and AMA loss to address the domain gap between text-to-2D and text-to-3D generation, our proposed framework exhibits substantial advancements over prior methods in text-to-3D generation. In summary, our study contributes to the field in three key aspects:
@@
  • To enhance the alignment between the generated results and the camera perspective, we propose CG-LoRA, which leverages camera information to dynamically generate CG-LoRA parameters for 2D diffusion models.
@@
2. Related Work
2.1. Text-to-3D Content Creation
@@
In recent years, there has been a significant surge in interest surrounding the evolution of text-to-3D generation (Michel et al., 2022; Poole et al., 2022; Li et al., 2023a; Chen et al., 2023d; Li et al., 2024a; Sella et al., 2023; Wu et al., 2024; Yang et al., 2023; Fei et al., 2024b; Babu et al., 2023; Fei et al., 2024d). This growing field has been propelled, in part, by advancements in pretrained vision-and-language models, such as CLIP (Radford et al., 2021), as well as diffusion models like Stable Diffusion (Rombach et al., 2022) and Imagen (Saharia et al., 2022). Contemporary text-to-3D models can generally be classified into two distinct categories: the CLIP-based text-to-3D approach and the diffusion-based text-to-3D approach. The CLIP-based text-to-3D approach (Michel et al., 2022; Ma et al., 2023b; Lei et al., 2022; Jain et al., 2022; Mohammad Khalid et al., 2022; Wei et al., 2023) employs CLIP encoders (Radford et al., 2021) to project textual descriptions and rendered images derived from the 3D object into a modal-shared feature space. Subsequently, CLIP loss (Jiang et al., 2023a; Wang et al., 2023b; Ding et al., 2023) is harnessed to align features from both modalities, optimizing the 3D representation to conform to the textual description. Various scholars have made significant contributions to this field. For instance, Michel et al. (Michel et al., 2022) are pioneers in proposing the use of CLIP loss to harmonize the text prompt with the rendered images of the 3D object, thereby enhancing text-to-3D generation. Ma et al. (Ma et al., 2023b) introduce dynamic textual guidance during 3D object synthesis to improve convergence speed and generation performance. However, these approaches have inherent limitations, as they tend to generate 3D representations with a scarcity of geometry and appearance detail. To overcome this shortcoming, the diffusion-based text-to-3D approach (Poole et al., 2022; Lin et al., 2023; Chen et al., 2023a; Tang et al., 2023; Jiang et al., 2023b; Huang et al., 2023a) leverages pretrained text-to-2D diffusion models (Rombach et al., 2022; Saharia et al., 2022) to guide the optimization of 3D representations. Central to these models (Cao et al., 2023; Richardson et al., 2023; Huang et al., 2024; Tang et al., 2023; Li et al., 2023b) is the application of SDS loss (Poole et al., 2022) to align the rendered images stemming from a variety of camera perspectives with the textual description. Specifically, given the target text prompt, Lin et al. (Lin et al., 2023) leverage a coarse-to-fine pipeline to generate high-resolution 3D content. Chen et al. (Chen et al., 2023a) decouple geometry modeling and appearance modeling to generate realistic 3D assets. For specific purposes, some researchers (Seo et al., 2023; Wang et al., 2023c) integrate trainable LoRA (Hu et al., 2021) branches into pretrained diffusion models. For instance, Seo et al. (Seo et al., 2023) put forth 3DFuse, a model that harnesses the power of LoRA to comprehend object semantics. Wang et al. (Wang et al., 2023c) introduce ProlificDreamer, where the role of LoRA is to evaluate the score of the variational distribution for 3D parameters. However, the LoRA parameter begins its journey from random initialization and maintains its independence from the camera and text. To address these limitations, we present two innovative modules: CG-LoRA and AMA loss. These modules are designed to enhance the model’s ability to consider important camera parameters and prioritize the generation of foreground objects throughout the text-to-3D creation process.
Figure 3. Overview of the proposed X-Dreamer, which consists of geometry learning and appearance learning.
2.2. Low-Rank Adaptation (LoRA)
@@
Low-Rank Adaptation (LoRA) (Hu et al., 2022) is a technique used to reduce memory requirements when fine-tuning a large model (Wang et al., 2024; Achiam et al., 2023; Li et al., 2024b; Chen et al., 2023e; Liu et al., 2024; Luo et al., 2024; Li et al., 2023c). It involves injecting only a small set of trainable parameters into the pretrained model, while keeping the original parameters fixed. During the optimization process, gradients are passed through the fixed pretrained model weights to the LoRA adapter, which is then updated to optimize the loss function. LoRA has been applied in various fields, including natural language processing (Hu et al., 2021; Brown et al., 2020), image synthesis (Zhang et al., 2023) and 3D generation (Seo et al., 2023; Wang et al., 2023c). To achieve low-rank adaptation, a linear projection with a pretrained weight matrix $\mathbf{W}_{0}\in\mathbb{R}^{d_{in}\times d_{out}}$ is augmented with an additional low-rank factorized projection. This augmentation is represented as $\mathbf{W}_{0}+\Delta\mathbf{W}=\mathbf{W}_{0}+\mathbf{A}\mathbf{B}$, where $\mathbf{A}\in\mathbb{R}^{d_{in}\times r}$, $\mathbf{B}\in\mathbb{R}^{r\times d_{out}}$, and $r\ll\min(d_{in},d_{out})$. During training, $\mathbf{W}_{0}$ remains fixed, while $\mathbf{A}$ and $\mathbf{B}$ are trainable. The modified forward pass, given the original forward pass $\mathbf{Y}=\mathbf{X}\mathbf{W}_{0}$, can be formulated as follows:
(1) $\mathbf{Y}=\mathbf{X}\mathbf{W}_{0}+\mathbf{X}\mathbf{A}\mathbf{B}.$
@@
3. Approach
3.1. Architecture
@@
In this section, we present a comprehensive introduction to the proposed X-Dreamer, which consists of two main stages: geometry learning and appearance learning. For geometry learning, we employ DMTet (Shen et al., 2021) as the 3D representation. DMTet is an MLP parameterized with $\Phi_{dmt}$ and is initialized with a 3D ellipsoid using the mean squared error (MSE) loss $\mathcal{L}_{MSE}$. Subsequently, we optimize DMTet and CG-LoRA using the SDS loss (Poole et al., 2022) $\mathcal{L}_{SDS}$ and the proposed AMA loss $\mathcal{L}_{AMA}$ to ensure the alignment between the 3D representation and the input text prompt. For appearance learning, we leverage bidirectional reflectance distribution function (BRDF) modeling (Torrance and Sparrow, 1967) following the previous approach (Chen et al., 2023a). Specifically, we utilize an MLP with trainable parameters $\Phi_{mat}$ to predict surface materials. Similar to the geometry learning stage, we optimize $\Phi_{mat}$ and CG-LoRA using the SDS loss $\mathcal{L}_{SDS}$ and the AMA loss $\mathcal{L}_{AMA}$ to achieve alignment between the 3D representation and the text prompt. Fig. 3 provides a detailed depiction of our proposed X-Dreamer.
3.1.1. Geometry Learning
@@
For geometry learning, an MLP network $\Phi_{dmt}$ is utilized to parameterize DMTet as a 3D representation. To enhance the stability of geometry modeling, we employ a 3D ellipsoid as the initial configuration for DMTet $\Phi_{dmt}$. For each vertex $v_{i}\in V_{T}$ belonging to the tetrahedral grid $T$, we train $\Phi_{dmt}$ to predict two important values: the SDF value $s(v_{i})$ and the deformation offset $\delta(v_{i})$. To initialize $\Phi_{dmt}$ with the 3D ellipsoid, we sample a set of $N$ points $\{p_{i}\in\mathbb{R}^{3}\}|_{i=1}^{N}$ approximately distributed on the surface of an ellipsoid and compute the corresponding SDF values $\{{SDF}(p_{i})\}|_{i=1}^{N}$. Subsequently, we optimize $\Phi_{dmt}$ using MSE loss. This optimization process ensures that $\Phi_{dmt}$ effectively initializes DMTet to resemble the 3D ellipsoid. The formulation of the MSE loss is given by:
(2) $\mathcal{L}_{MSE}=\frac{1}{N}\sum_{i=1}^{N}\left(s(p_{i};\Phi_{dmt})-SDF(p_{i})\right)^{2}.$
@@
After initializing the geometry, our objective is to align the geometry of DMTet with the input text prompt. Specifically, we generate the normal map $\boldsymbol{n}$ and the object mask $\boldsymbol{m}$ from the initialized DMTet $\Phi_{dmt}$ by employing a differentiable rendering technique (Torrance and Sparrow, 1967), given a randomly sampled camera pose $\boldsymbol{c}$. Subsequently, we input the normal map $\boldsymbol{n}$ into the frozen stable diffusion (SD) with a trainable CG-LoRA and update $\Phi_{dmt}$ using the SDS loss, which is defined as follows:
@@
(3) $\small\nabla_{\Phi_{dmt}}\mathcal{L}_{\mathrm{SDS}}=\mathbb{E}_{t,\boldsymbol{\epsilon}}\left[w(t)\left(\hat{\boldsymbol{\epsilon}}_{\Theta}(\boldsymbol{n}_{t};\boldsymbol{y},t)-\boldsymbol{\epsilon}\right)\frac{\partial\boldsymbol{n}}{\partial{\Phi_{dmt}}}\right],$
@@
where $\Theta$ represents the parameter of SD, $\hat{\boldsymbol{\epsilon}}_{\Theta}(\boldsymbol{n}_{t};\boldsymbol{y},t)$ denotes the predicted noise of SD given the noise level $t$ and text embedding $\boldsymbol{y}$. Additionally, $\boldsymbol{n}_{t}=\alpha_{t}\boldsymbol{n}+\sigma_{t}\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\boldsymbol{I})$ represents noise sampled from a normal distribution. The implementation of $w(t)$, $\alpha_{t}$, and $\sigma_{t}$ is based on the DreamFusion (Poole et al., 2022).
@@
(4) $\mathcal{L}_{AMA}=\frac{1}{L}\sum_{i=1}^{L}|\boldsymbol{a}_{i}-\eta(\boldsymbol{m})|,$
@@
Figure 4. Illustration of Camera-Guided Low-Rank Adaptation (CG-LoRA).
3.1.2. Appearance Learning
After obtaining the geometry of the 3D object, our objective is to compute its appearance using the Physically-Based Rendering (PBR) material model (McAuley et al., 2012). The material model comprises the diffuse term $\boldsymbol{k}_{d}\in\mathbb{R}^{3}$, the roughness and metallic term $\boldsymbol{k}_{rm}\in\mathbb{R}^{2}$, and the normal variation term $\boldsymbol{k}_{n}\in\mathbb{R}^{3}$. Firstly, the Diffuse Term $\boldsymbol{k}_{d}\in\mathbb{R}^{3}$ accurately represents how the material responds to diffuse lighting. This term is captured by a three-dimensional vector that denotes the material’s diffuse color. Secondly, the Roughness and Metallic Term $\boldsymbol{k}_{rm}\in\mathbb{R}^{2}$ effectively captures the material’s roughness and metallic properties. It is represented by a two-dimensional vector that conveys the values of roughness and metallicness. Roughness quantifies the surface smoothness, while metallicness indicates the presence of metallic properties in the material. These terms collectively contribute to the overall visual realism of the rendered scene. Additionally, the Normal Variation Term $\boldsymbol{k}_{n}\in\mathbb{R}^{3}$ plays a crucial role in characterizing variations in surface normals. This term, typically used in conjunction with a normal map, enables the simulation of fine details and textures on the material’s surface. During the rendering process, these three PBR material terms interact to accurately simulate the optical properties exhibited by real-world materials. This integration facilitates rendering engines in achieving high-quality rendering results and generating realistic virtual scenes.
For any point $p\in\mathbb{R}^{3}$ on the surface of the geometry, we utilize an MLP parameterized by $\Phi_{mat}$ to obtain the three material terms, which can be expressed as follows:
(5) $(\boldsymbol{k}_{d},\boldsymbol{k}_{n},\boldsymbol{k}_{rm})=\text{MLP}\left(\mathcal{P}(p);\Phi_{mat}\right),$